If P(A or B) = 0.8, P(A) = 0.8 and P(B) = 0.5, determine P(A and B).
(Points : 3)
0.35
0.50
0.64
0.13

Answer by robertb(4012)

P(A and B) = P(A) + P(B) - P(A U B) = 0.8 + 0.5 - .8 = 0.5